5. Solve the initial value problem y” + 2y + y = 0, y(0=1, y(1)=0 (a) y=e* +4xe* (b) y= e^ +3xe 1 (c) y= e^ +2xe (d) y=
Answer #1

Given differential equation is y+2y+ y = 0 (D*+2D+1)y=0 where De Auxiliary equation is ma + 2m +1 = 0 (m+1) = 0 m=-1,-1 ::